# FacAID: A Transformer Model for Neuro-Symbolic Facade Reconstruction

### v1

@@ 7. Model Ablations
We have trained our model on 5 different segmentation discretization resolutions. Starting from a fine grid—1000x1000, all the way to a very coarse one—50x50. The results (cf Figure 7(b)) show that the resolution has little to no impact on the model’s ability to learn. We assume that this might be caused by the fact that the model learns the structure mostly based on the order and counts of the rectangles that form the input segmentation and their geometric parameters (positions and sizes) just offer additional guidance in which case even the coarse discretization grid should be sufficient. After all, for a single facade structure, there exist multiple sizing variations.
@@
Since the model is composed of two distinct elements—encoder and decoder—we have tested whether making one of them bigger than the other would be beneficial to the learning process. We observe that the model with a deeper decoder tends to overfit a bit faster while making the encoder deeper produces the opposite result (cf Figure 7(e)). That is why we opt for a balanced architecture, as the validation curve of such a model correlates most closely with its training counterpart.
@@
The model was also tested against its overall width and depth. We have varied those parameters by a factor of 2 for each of them. The results show that the model performs best when the number of its embedding dimensions (width) greatly surpasses its number of layers (cf Figure 7(d)). We conclude that this result most likely comes from the fact that the transformer model architecture already provides a lot of depth in its structure by feeding the embeddings through the quite shallow blocks multiple times. This allows the embeddings to pick up additional context information while their larger size provides the space required to store more of the nuances of the segmentation structure.
@@
The data that we use for training is procedurally generated by our data generation module. This allows us to generate as many facades as we require from the facade data distribution that the data generator represents. We use this fact to test how the amount of data the our model is trained on influences its ability to reconstruct facades. We generate 4 datasets of varying sizes—starting from 20k facades all the way up to 200k facades. Since the space of all possible facade segmentations is vast, we do not expect any data-point repeats, even in the largest dataset. Our experiments show that the dataset size has a big impact on the abilities of the model—the more data it was fed the better the results (cf. Figure 7(a)). We assume that this behavior is caused by the complexity of the function being learned which arises from subspaces, in the all possible segmentations space, that should evaluate to the same structure even though their sizings (and by consequence the sequence that represents them) differ vastly.
@@
We have performed multiple tests to identify what values of learning rate and batch size best fit our model. From the results (cf. Figures 7(f) and 7(g)) we conclude that learning rates equal to 0.0001 and 0.00001 both provide satisfactory results (we opt for the smaller value), while that batch size does not impact the result substantially as long as it’s not too big (we settled on a batch size of 32).
@@
8. Qualitative Results
The qualitative results were generated on facades manually segmented into grids of rectangular boxes, and annotated. The segmentations have been crafted such that they match the terminals of our grammar and take around 5 minutes each to construct. For that task we have implemented a rudimentary segmentation interface. In practice, our goal would be to either use automatic segmentation tools, based on machine learning methods (Selvaraju et al., 2021) or on interactive segmentation tools (Musialski et al., 2012). However, generally, we consider that as given input which is not part of and essentially does not influence the contribution of this paper.
@@ 9. Discussion and Limitations
While our method reproduces a vast range of facades, it has limitations tied to the split-grammar language it relies on. If no derivation trees based on the grammar can reproduce a given facade, the model may not deliver a correct result (Figure 8). This can be mitigated by extending the grammar with more procedures and making them more general, which requires insight into architectural design. Additionally, the transformer model’s memory requirement scales proportionally to the sequence length squared, limiting facade sizes and detail levels.

### v2

@@ 8. Model Ablations
We have trained our model on 5 different segmentation discretization resolutions. Starting from a fine grid—1000x1000, all the way to a very coarse one—50x50. The results (cf. Figure 7(b)) show that the resolution has little to no impact on the model’s ability to learn. We assume that this might be caused by the fact that the model learns the structure mostly based on the order and counts of the rectangles that form the input segmentation and their geometric parameters (positions and sizes) just offer additional guidance in which case even the coarse discretization grid should be sufficient. After all, for a single facade structure, there exist multiple sizing variations.
@@
Since the model is composed of two distinct elements—encoder and decoder—we have tested whether making one of them bigger than the other would be beneficial to the learning process. We observe that the model with a deeper decoder tends to overfit a bit faster while making the encoder deeper produces the opposite result (cf. Figure 7(e)). That is why we opt for a balanced architecture, as the validation curve of such a model correlates most closely with its training counterpart.
@@
The model was also tested against its overall width and depth. We have varied those parameters by a factor of 2 for each of them. The results show that the model performs best when the number of its embedding dimensions (width) greatly surpasses its number of layers (cf. Figure 7(d)). We conclude that this result most likely comes from the fact that the transformer model architecture already provides a lot of depth by feeding the embeddings through the quite shallow blocks multiple times. This allows the embeddings to pick up additional context information while their larger size provides the space required to store more of the nuances of the segmentation structure.
@@
The data that we use for training is procedurally generated by our data generation module. This allows us to generate as many facades as we require from the facade data distribution that the data generator represents. We use this fact to test how the amount of data our model is trained on influences its ability to reconstruct facades. We generate 4 datasets of varying sizes—starting from 20k facades all the way up to 200k facades. Since the space of all possible facade segmentations is vast, we do not expect any data-point repeats, even in the largest dataset. Our experiments show that the dataset size has a big impact on the abilities of the model—the more data it was fed, the better the results (cf. Figure 7(a)). We assume that this behavior is caused by the complexity of the function being learned which arises from subspaces, in the all possible segmentations space, that should evaluate to the same structure even though their sizings (and by consequence the sequence that represents them) differ vastly.
@@
We have performed multiple tests to identify what values of learning rate and batch size best fit our model. From the results (cf. Figures 7(f) and 7(g)) we conclude that learning rates equal to 0.0001 and 0.00001 both provide satisfactory results (we opt for the smaller value), while the batch size does not impact the result substantially as long as it’s not too big (we settled on a batch size of 32).
@@ 9. Discussion and Limitations
While our method reproduces a vast range of facades, it has limitations tied to the split-grammar language. If no derivation tree based on the grammar can reproduce a given facade, the model will not deliver a correct result (Figure 9). This can be mitigated by extending the grammar with more procedures and making them more general, which requires insight into architectural design. Additionally, the transformer model’s memory requirement scales proportionally to the sequence length squared, limiting facade sizes and detail levels.
@@ 10. Conclusions and Future Work
Figure 10. Results of our method: Each sequence of 5 images first depicts the original target image and its segmentation (partially from (Musialski et al., 2012)). Some of the inputs are purposefully noisy to showcase the robustness of our approach. The first image after each vertical line showcases the procedural reconstruction generated by our method, and the last two results are procedurally generated variations. Best viewed in close-up in the electronic version.
Figure 11. Rendering of the results of our method including material properties assigned to different segmentation classes. Each sequence shows the input image, the procedural reconstruction, and two variations of the same facade procedure. Best viewed in close-up in the electronic version.